Evaluate the function given below for the following values: (a) x = –2, (b) x = 8.
f(X)=20x-4

Answer by fractalier(6550)

To evaluate a function at a given value, merely plug in the number(s) given for x...as in our function
f(x) = 20x - 4
we get
f(-2) = 20(-2) - 4 = -40 - 4 = -44
f(8) = 20(8) - 4 = 160 - 4 = 156